OpenStudy (anonymous):

Please help me with this question The regular price of a child's entry ticket to a water park is $6 less than that for an adult's. The park offers half off all entry tickets during the off-peak season. The Jordans paid a total of $69 for 2 adult tickets and 1 child's ticket to the water park during the off-peak season. The following equation represents this situation, where x represents the regular price of an adult ticket. : 132 = one-half x + 2(x - 9) What is the regular price of a child's ticket? @e.mccormick

OpenStudy (anonymous):

@StudyGurl14

OpenStudy (anonymous):

@e.mccormick doesn't help, I waited for him lik 40 min and didnt even look at my question, he is looking at others but he will never help

OpenStudy (chosenmatt):

first multiply both sides of the equation by 2 then you solve for X :)

OpenStudy (anonymous):

I multiplied both sides by two but I'm still having trouble solving for x

OpenStudy (anonymous):

could you help some more?

OpenStudy (anonymous):

@abb0t

OpenStudy (chosenmatt):

okay one sec

OpenStudy (anonymous):

@StudyGurl14

OpenStudy (studygurl14):

Okay one sec while i type it

OpenStudy (studygurl14):

\(\large 132=\frac{1}{2}x+2(x-9)\) First step is to distribute the 2 to the x and -9 \(\large 132=\frac{1}{2}x+2x - 18\)

OpenStudy (studygurl14):

Next step. Move the -18 to teh other side by adding it to both sides. It cancels out on the right side. What is 132 + 18?

OpenStudy (anonymous):

150

OpenStudy (studygurl14):

Right. So you're left with \(\large 150=\frac{1}{2}x+2x\)

OpenStudy (studygurl14):

1/2 = 0.5 What is 2 + 0.5 ?

OpenStudy (anonymous):

2.5

OpenStudy (anonymous):

why add 2? @StudyGurl14

OpenStudy (studygurl14):

So you get... \(\large 150=2.5x\)

OpenStudy (studygurl14):

You add 2 because you have 2x

OpenStudy (anonymous):

ok

OpenStudy (studygurl14):

Next step: Divide 150 by 2.5 and you'll get the answer

OpenStudy (anonymous):

I got 60 but there is no option on the answer list for 60

OpenStudy (anonymous):

These are the possible answers: $42 $48 $66

OpenStudy (anonymous):

@StudyGurl14

OpenStudy (studygurl14):

ok, that's odd. hold on whie i check something...

OpenStudy (studygurl14):

Ok @M23454 Was it: \(\large\frac{1}{2x}\) or \(\large\frac{1}{2}x\) ??

OpenStudy (anonymous):

it was 1/2 x

OpenStudy (studygurl14):

then...i don't know...it should be 60...doesn't make sense

OpenStudy (studygurl14):

u sure you typed the problem right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

I just got the answer, it was 48

OpenStudy (anonymous):

thanks for your help!

OpenStudy (anonymous):

maybe I just confused you the way I told you the problem